(don't know if I have chosen the correct topic, but please help me.) How do I select the value of x that are solutions to F(x)=0, when F(x)=-5x+9x+3
Answer by rothauserc(4718)

we are given,
0 = -5x +9x +3 then
4x + 3 = 0
x = -3/4
check the answer
0 = -5(-3/4) +9(-3/4) + 12/4
0 = 15/4 -27/4 +12/4
0 = 27/4 -27/4
0 = 0
our answer checks